OpenStudy (anonymous):

how many solutions 4x-14y=6 -2x+7y=-3

OpenStudy (anonymous):

whole solutions?

OpenStudy (anonymous):

how many solutions are there to the following system of equations? sorry

OpenStudy (anonymous):

4x-14y=6 add 14y to both sides and subtract 6 from both sides. 4x-6-14y+14y=6-6+14y 4x-6=14y divide both sides by 14 \[y=\frac{ 4x-6 }{ 14 }=\frac{2(2x-3)}{14}=\frac{2x-3}{7}\]

OpenStudy (anonymous):

look. the first equation is the line. There are infinetely many solutions.

OpenStudy (anonymous):

for any x there will be one y

OpenStudy (anonymous):

the same thing is with second equation.